OpenStudy (anonymous):

Solve the equation. Show your work. 8x – 12 = 4x + 24

OpenStudy (anonymous):

8x – 12 = 4x + 24 +12 +12 8x =4x +24 +12 4x=36 x=36/4 =9

OpenStudy (anonymous):

You got it.